OpenStudy (anonymous):

Algebra 2 help?? Multiply and simplify if possible. ^4sqrt11 x ^4sqrt10

OpenStudy (anonymous):

\[\sqrt[4]{11}\times \sqrt[4]{10}\]

OpenStudy (anonymous):

okay this is easiest on a calculator because neither number can be fourth rooted properly. but essentially when you fourth root something like you are trying to do here, you look for a number that when multiplied by itself four times it will get whatever number is inside the little rooting box. :)

OpenStudy (marissalovescats):

Well no, you wouldn't want to do it on a calculator for that reason. Also notice it asks you to simplify if you can. You will notice both are fourth roots so what you want to do is do this: